Write the slope-intercept equation for the line that passes through (-7, 6) and is perpendicular to -7x + 9y = -2 Please show all of your work.
Answer by mananth(16946)

-7x-9y=-2
Find the slope of this line
9y=-7x+2
Divide by 9
y=- 7/9 x 0.22
Compare this equation with y=mx+b
slope m =-0.78

The slope of a line perpendicular to the above line will be the negative reciprocal -0.78
m1*m2=-1
The slope of the required line will be 1 2/7, (1 2/7=1.29)

m=1.29 ,point(4 ,-2)
Find b by plugging the values of m & the point in
y=mx+b
-2=5.14+ b
b=-7.14
m=1.29
The required equation is y=1.29x-7.14